4,-1,5,3,-2 4,-1,5,3,-2 4,-1,5,3,-2.....
find the sum of the first 89 terms of the sequence... I don't need this solved but can you show me an example please

Answer by Theo(13342)

This looks like a repeating sequence of numbers.

You have 4, -1, 5, 3, -2
4, -1, 5, 3, -2
4, -1, 5, 3, -2
.....

You want the sum of the first 89 terms in the sequence.

89 / 5 = 17.8

.8 * 5/5 = 4/5

17.8 is equivalent to 17 + 4/5

You're going to go through the sequence completely 17 times and then you will end up on the 4th term of the last sequence.

The sum of this sequence would be calculated as follows:

The sum of all 5 numbers in each sequence is equal to 9.

the sum of the first 4 numbers in the last sequence is equal to 11.

17 * 9 = 153

Add the first 4 numbers in the last sequence to that and you get:

153 + 11 = 164.
The sum of the first 89 terms in the sequence is equal to 164.